# Multilayer Network of Cardiovascular Diseases and Depression via Multipartite Projection

### v1

@@ Abstract
There is a significant comorbidity between cardiovascular diseases (CVD) and depression that is highly predictive of poor clinical outcome. Yet, its underlying biological pathways remain challenging to decipher, presumably due to its non-linear associations across multiple mechanisms. Mutual information provides a framework to analyze such intricacies. In this study, we proposed a multipartite projection method based on mutual information correlations to construct multilayer disease networks. We applied the method to a cross-sectional dataset from a wave of the Young Finns Study. This dataset assesses CVD and depression, along with related risk factors and two omics of biomarkers: metabolites and lipids. Instead of directly correlating CVD-related phenotypes and depressive symptoms, we extended the notion of bipartite networks to create a multipartite network that connects these phenotype and symptom variables to intermediate biological variables. Projecting from these intermediate variables results in a weighted multilayer network, where each link between CVD and depression variables is marked by its ‘layer’ (i.e., metabolome or lipidome). Using this projection method, we identified potential mediating biomarkers that connect CVD to depression. These included metabolites such as creatinine, valine, phospholipids in very large high-density lipoproteins HDL, triglycerides in small low-density lipoproteins (LDL), and apolipoprotein B; and lipids such as specific sphingomyelins (SM), phosphatidylcholines (PC), triacylglycerols (TAG) and diacylglycerols (DAG). These biomarkers thus may play significant roles in the biological pathways of CVD-depression comorbidity. Additionally, the projected network highlights sex and BMI as the most important risk factors, or confounders, associated with the comorbidity. Our method can generalize to any number of omics layers and disease phenotypes, offering a truly system-level overview of biological pathways contributing to comorbidity.
@@ Introduction
Cardiovascular diseases (CVD) and depression are pervasive health challenges. The significant co-occurrence of these conditions indicates a profound comorbidity [1, 2, 3]. Indeed, individuals with depression face an elevated risk of developing CVD compared to the general population [4]. Conversely, individuals with CVD are more prone to developing depression, and their comorbidity is linked to a 3-fold higher mortality rate [5, 6]. In either direction, both behavioral and biological factors play a role [5, 7, 8]. An extensive literature documents possible biological mechanisms underlying the bidirectional relationship between CVD and depression, such as increased inflammation, endocrine disturbances, and imbalances in neurotransmitter levels, sympathetic and parasympathetic nervous systems [9, 10, 11, 12]. Either phenotype alone already significantly diminishes the quality of life, imposes a considerable burden on the healthcare system, and remains a challenge to treat —their combination presents an even bigger challenge. Success in solving this major public health concern, e.g., developing effective preventive and treatment strategies, hinges on accurately identifying potential pathways that connect CVD and depression. However, the precise nature of these interactions has not yet clearly been established. Continued effort to understanding their biological associations, by identifying molecular biomarkers and uncovering underlying risk factors, remains important. In this article, we contribute to this effort by introducing a novel disease network construction method, investigating the relationships between CVD-related phenotypes and depressive symptoms, and their biological mediators.
@@
The concept of ’human disease networks’ has been introduced for this purpose [13], providing a powerful framework to study human diseases and underlying biological mechanisms using a network perspective. As an illustrative example, Goh et al. introduced a projection method to reconstruct disease networks based on a bipartite graph [13]. In their projected networks, the association between two diseases was established if there was a gene that was implicated in both. The strength of gene-disease associations was quantified by (linear) Pearson correlation, and disease-disease associations (i.e., explaining their comorbidity) were defined as the number of genes that were significantly associated with both diseases. However, there were several limitations of this methodology that might lead to bias in estimating disease associations. Firstly, it used only genes to connect diseases, whereas in reality there are multiple omics (e.g., genes, proteins, lipids, etc.) connecting diseases. Secondly, simply counting the number of shared intermediate biomarkers was a coarse measure that treated each biomarker equally, no matter their strength of (Pearson) correlation. Thirdly, the use of Pearson correlation means that potentially nonlinear correlations were missed [14]. Lastly, these methods did not allow for the incorporation of higher-order correlations into the projection.
@@
Some studies have expanded on the methodology to address the bias. Park, J. et al. implemented cellular networks and incorporated genes and proteins into their analyses  [15]. To tackle the second limitation, Zhou, X. et al. applied the cosine similarity of respective symptom vectors of given diseases to quantify disease associations, and defined the order of interactions based on the length of path between proteins [16]. Grosdidier, S. et al. proposed a concept of molecular comorbidity index (MCI) based on proteins common between two diseases to estimate the strength of the association [17]. Although extending to multiple omics or defining a more complicated disease association, methodological shortcomings remain. Specifically, none of the current methods provide a combined solution that incorporates multiple omics datasets into a single network analysis, integrates a more refined counting of mediating biomarkers, and utilizes an expressive non-linear correlation.
@@
To fill the above gap, we propose a multipartite projection method that (i) can combine any number of omics datasets and diseases; (ii) is based on the non-parametric Shannon mutual information (MI) correlation; and (iii) counts shared biomarkers and risk factors in a weighted manner. MI is a non-parametric metric that captures both linear and nonlinear mutual interdependence between two variables [18]. Counting biomarkers in a weighted way provides a finer-grained measure to quantify the links between diseases. For this purpose, we propose three definitions for the projected disease links. These definitions not only count the number of their shared neighboring (intermediate) biomarkers, but also consider the strength of their nonlinear correlations with these biomarkers, the importance of these correlation links, and their own neighboring biomarkers that have the shortest path. Compared to direct correlation, the links implied by the network projection provides deeper insights into potential common etiologies.
@@
We applied the projection method to a dataset from the Young Finns Study (YFS) and selected a projection definition that performs better. The dataset encompasses three sets of variables: (1) CVD and depression phenotypes (see Table 1 and  2) and related risk factors (e.g., covariates, Table 4), (2) metabolites, and (3) lipids. Our method and its resulting weighted network allow us to: (i) in parallel identify important biomarkers, representing multiple biological layers possibly related to the development of CVD-depression comorbidity; (ii) within this network, integrate and account for important risk factors; (iii) identify biomarkers related to specific risk-disease and disease-disease associations. The projection method detects sex and body mass index (BMI) as two important risk factors related to both CVD and depression, and identifies significant mediating biomarkers underlying their comorbidity, including metabolites such as creatinine, valine, phospholipids in very large high-density lipoproteins HDL, triglycerides in small low-density lipoproteins (LDL), and apolipoprotein B, as well as lipids such as specific sphingomyelins (SM), phosphatidylcholines (PC), triacylglycerols (TAG) and diacylglycerols (DAG). By including any number of omics datasets and taking into account non-linear correlations, our method offers a truly system-level overview of biological pathways contributing to a comorbidity.
@@ Methods
Our proposed method was applied to the 2007 wave of the Young Finns Study (YFS) dataset. This cross-sectional follow-up study presented a multidimensional array of data that encompasses (1) CVD-related phenotypes, depressive symptoms, and related risk factors, (2) metabolomic data, and (3) lipidomic data. This multidimensional dataset enables the identification of biomarkers, and the assessment of the effects of risk factors on both CVD and depression. Notably, it contains two omics datasets, allowing for multipartite projection to calculate projected scores between diseases.
For each participant, phenotype variables consist of 17 CVD-related indicators and six risk factors (see Table 1 and  4). Depressive variables include Beck’s score, and 21 individual depressive symptoms that contribute to the calculation of the summary score (see Table 2). The two omics datasets comprise 228 metabolomic variables and 437 lipidomic variables saved in two separate tables.
@@
We used random sample imputation to replace each missing value of a variable with a random sample drawn from existing values of that variable. This is conservative in the sense that it cannot introduce spurious correlations, and the more missing values a variable has, the lower its correlations. Following that, we retained the data points in which data of all omics variables were available, resulting in a total of 1686 participants (983 women, 703 men). The continuous variables were discretized by binning the data using a quantile-based method, dividing each continuous variable into probabilistically equal-sized buckets. Discretization is needed because part of the data is discrete and information-theoretic metrics cannot be compared for mixed data. The optimal number of quantiles was determined by Sturges’ rule. Categorical and discrete variables are included unchanged.
@@
Before constructing the significant MI network, we first applied a filtering process to minimize double counting and reduce redundant information among variables. The step was necessary because our proposed method incorporates the strength of MI correlations into the projection, and ensuring that each correlation is uniquely represented without redundancy is crucial for the method’s effectiveness. Consequently, some repeated variables and one variable from each highly correlated pair based on MI correlation were excluded from the analysis (see S1.1 in the Supplementary Information). This included Beck’s score, CVD-related variables (such as "fmd40", "fmd40pr", "fmd60", "fmd60pr", "fmd80", "fmd80pr", "maxfmd", "imtmax", "bbmax" and "volscore"), and another 115 variables in metabolites and lipids. This process resulted in a final dataset of 584 variables for further analysis.
@@
The MI can quantify correlations directly between variables of phenotypes, symptoms and risk factors. However, these direct correlations would not be able to suggest any plausible causative mechanisms (pathways) between them, let alone identify key biomarkers. Our multipartite projection method computes these correlations indirectly as projected scores, based on their MI correlations with intermediate biomarkers.
@@
In the multipartite projection method, we defined the projected score as the sum of the average MI correlations between each pair of variables and their shared neighboring nodes (intermediate biomarkers). The definition is formulated as follows:
@@
This projection definition considers intermediate biomarkers in a weighted manner. When projected to different types of biomarkers this leads to a weighted, multilayer disease network of CVD phenotypes, depressive symptoms and related risk factors, in which each layer of the network corresponds to the type of biomarker involved in the projection. A stylized representation of this projection definition using average MI correlation is depicted in Figure 1. We also explored alternative definitions of the projected score, and selected this particular definition based on a comparative analysis (see Section S2.1 in the Supplementary Information).
@@
The projection method allows us to calculate the contribution of each biomarker (in our case, metabolites and lipids) to each projected score. Considering the projection definition used in the main analysis and our focus on CVD-depression comorbidity, we defined the total contribution score of a biomarker to all possible projected links between CVD phenotypes and depressive symptoms as follows:
@@
where $Y_{k}$ is a biomarker, $X_{i}$ and $X_{j}$ represent pairs of phenotypes and symptoms. Both $f(X_{i};Y_{k})$ and $f(X_{j};Y_{k})$ are non-zero. Such a contribution score enables the identification of significant biomarkers that contribute to the overall comorbidity between CVD and depression. For a given pair of $(X_{i},X_{j})$, the method identifies the significant biomarkers that contribute to the specific projected score.
@@
It is widely recognized that CVD and depression are often linked to risk factors [19, 12]. When risk factors are included in the projection, the projected network provides a means of assessing the relative importance of risk factors in relation to a specific phenotype. The relative importance of a risk factor with respect to a particular phenotype is defined as the ratio of the projected score between the phenotype and the risk factor to the total projected score between the phenotype and all risk factors. This can be formulated as:
@@
where $X_{j}$ is a phenotype and $Z_{i}$ is one of the risk factors that has a link to the phenotype. $w(Z_{i};X_{j})$ is the projected score between the phenotype and the risk factor. The relative importance of risk factors considering CVD and depression is computed as the average of $r(Z_{i},X_{j})$ over all CVD-related phenotype and depressive symptoms, respectively. For a particular strong risk-disease link, we can identify the significant biomarkers associated with the relationship using Equation  3.
@@ Results
Significant tripartite MI correlation network and projected multilayer networks
@@
The projection method builds on a significant multipartite MI correlation network. Figure 2A displays the multipartite (tripartite, in our case) network, in which variables are partitioned into three groups: metabolome (blue), lipidome (green), and the combined group of phenotypes/symptoms of CVD (yellow) and depression (red) along with risk factors (purple). This network is tripartite because it shows only the links between metabolites/lipids and variables in the combined group. The edge width is determined by the mean MI value, calculated over 20 runs of random sample imputation. The node size is proportional to its weighted degree. To enhance readability and visualization, nodes in the tripartite network were further filtered based on their weighted degree, retaining 105 nodes and 669 edges. No depressive symptom variables appear in this network, as these variables exhibit limited correlations with biomarker variables. It is important to note that this filtering process is only for the purpose of network visualization; the projection analysis was performed on the network before filtering. Therefore, the depressive symptoms were still included in the analysis.
@@
Observing the tripartite network, risk factors such as sex and BMI, CVD-related phenotypes like the ideal CVH score, average diastolic, and systolic blood pressure, exhibit a high weighted degree. This implies their extensive correlations with biomarker variables. However, based solely on this network, it is difficult to understand which biomarkers are driving these strong correlations, thereby obscuring the underlying causative mechanisms and biological pathways.
@@
Figure 2B and C illustrate the metabolomic layer and lipidomic layer of the projected network of CVD-related phenotypes and depressive symptoms, respectively. In these projected networks, the link (edge) width is quantified by the mean projected score over 20 runs of random sample imputation. The node size is proportional to its weighted degree. We only include the projected links between CVD and depression variables due to our focus on the comorbidity between these two conditions. The projected networks are thus displayed in a bipartite layout. In the metabolomic layer of the projected network, we find strong projected links between five individual depressive symptoms (pessimism (’b2’), guilty feelings (’b5’), crying (’b10’), changes in appetite (’b18’), loss of libido (’b21’)) and three CVD-related variables (average diastolic blood pressure (’dkv’), average systolic blood pressure (’syst’), and ideal CVD health score (’idealCVH’)). In the lipidomic layer, strong projected links are observed between three depressive symptoms (’b10’, ’b18’ and ’b21’) and the same three CVD-related phenotypes. These strong projected links suggest a high risk of co-occurrence between these CVD-related phenotypes and depressive symptoms, predicting multimorbid phenotypes of CVD and depression.
@@
To investigate the CVD-depression comorbidity, we calculated the mean total contribution score of each metabolite and lipid to all projected links between CVD-related phenotypes and depressive symptoms over 20 runs of random imputation. Figure 4A and B display the top ten mediating metabolites and lipids that contribute to these projected CVD-depression links. The top mediating metabolites include creatinine, valine, leucine, phospholipids in very large HDL and small VLDL, triglycerides in small LDL and very small VLDL, and apolipoprotein B. The top mediating lipids consist of several specific sphingomyelins (SM), phosphatidylcholines (PC), triacylglycerols (TAG), diacylglycerols (DAG), phosphatidylglycerols (PG), and phosphatidylethanolamines (PE). These metabolites and lipids were identified as significant biomarkers that link CVD and depression through biological pathways such as inflammation, cholesterol regulation, and atherosclerosis. Most of these metabolites and some lipids are already known, while others, such as SM and PG, are potentially novel biomarkers. Detailed rankings and further information on these biomarkers can be found in Table 3.
@@
Including risk factors in the projection results in a projected multilayer network, in which nodes are partitioned into three groups: risk factors, CVD-related phenotypes, and depressive symptoms (see Figure 5A and B). Observing both metabolomic and lipidomic layers of the projected network, the high weighted degree of sex and BMI indicates their extensive associations with both CVD and depression compared to other risk factors. The relative importance of all risk factors is shown in Table 4. The results suggest that sex and BMI are two important risk factors related to both CVD and depression, which corroborates previous studies [20, 21, 22, 23]. It is important to note that sex can be a covariate or confounder. Therefore, we conducted gender-specific analyses (see S2.3 in the Supplementary Information). Additionally, age–another well-known risk factor–does not show high importance in the YFS dataset, likely because the YFS participants were all young adults.
We took two demonstrated high-score projected links ("Sex"–"Appetite (’b18’)" and "Sex"–"idealCVH") as examples to further analyze the contributing biomarkers. In this case, sex may act as a confounding factor, influencing both variables. Therefore, it is plausible for appetite and the ideal CVH score to exhibit a strong projected score in the projected multilayer networks (see Figure 2). The significant mediating biomarkers based on the mean total contribution score are shown in Figure 5C, D. Details of these biomarkers can be found in Table S6 and S7. For the link between sex and appetite, significant metabolites include creatinine, free cholesterol in large HDL, and phospholipids, cholesterol and triglycerides in very large VLDL (very low-density lipoprotein). Significant lipids are mainly sphingomyelins (SM) 32:2 and lysophosphatidylcholines (LPC). For the link between sex and the ideal CVH score, significant metabolites encompass phospholipids, free and total cholesterol in large HDL and LDL, apolipoprotein B, and Leucine. Significant lipids are primarily triacylglycerols (TAG) and diacylglycerols (DAG). These identified biomarkers shed light on how risk factors influence the development of CVD and depression.

### v2

@@ Abstract
Cardiovascular diseases (CVD) and depression exhibit significant comorbidity, which is highly predictive of poor clinical outcomes. Yet, the underlying biological pathways remain challenging to decipher, presumably due to the non-linear associations across multiple mechanisms. In this study, we introduced a multipartite projection method based on mutual information correlations to construct multilayer disease networks as a novel approach to explore such intricate relationships. We applied this method to a cross-sectional dataset from a wave of the Young Finns Study, which includes data on CVD and depression, along with related risk factors and two omics of biomarkers: metabolites and lipids. Rather than directly correlating CVD-related phenotypes and depressive symptoms, we extended the notion of bipartite networks to create a multipartite network, linking these phenotypes and symptoms to intermediate biological variables. Projecting from these intermediate variables results in a weighted multilayer network, where each link between CVD and depression variables is marked by its ‘layer’ (i.e., metabolome or lipidome). Applying this projection method, we identified potential mediating biomarkers that connect CVD with depression. These included metabolites such as creatinine, valine, phospholipids in very large high-density lipoproteins (HDL), triglycerides in small low-density lipoproteins (LDL), and apolipoprotein B; as well as lipids such as specific sphingomyelins (SM), phosphatidylcholines (PC), triacylglycerols (TAG) and diacylglycerols (DAG). These biomarkers may therefore play significant roles in the biological pathways underlying CVD-depression comorbidity. Additionally, the network projection highlighted sex and BMI as key risk factors, or confounders, in this comorbidity. Our method is scalable to incorporate any number of omics layers and various disease phenotypes, offering a comprehensive, system-level perspective on the biological pathways contributing to comorbidity.
@@ Introduction
Cardiovascular diseases (CVD) and depression present major global health challenges with a significant co-occurrence yielding profound comorbidity [1, 2, 3]. Indeed, individuals with depression face an elevated risk of developing CVD [4]. Conversely, individuals with CVD are more prone to developing depression, and their comorbidity is linked to a 3-fold higher mortality rate [5, 6]. Thus, where each phenotype alone already significantly diminishes quality of life, imposes a considerable burden on the healthcare system, and remains challenging to treat —their combination presents an even bigger challenge. In either direction, both behavioral and biological factors play a role [5, 7, 8]. An extensive literature explored possible biological mechanisms underlying this bidirectional relationship, such as increased inflammation, endocrine disturbances, and imbalances in neurotransmitter levels, as well as dysregulation of the sympathetic and parasympathetic nervous systems [9, 10, 11, 12]. However, the precise nature of these interactions has not yet clearly been established and continued efforts to understand their biological associations remain important. In this article, we contribute to this effort by introducing a novel method for constructing disease networks, aimed at elucidating the relationships between CVD phenotypes, depressive symptoms, and their biological mediators.
@@
The concept of ’human disease networks’ has emerged as a powerful framework to study human diseases and underlying biological mechanisms using a network perspective [13]. As an illustrative example, Goh et al. introduced a projection method to reconstruct disease networks based on a bipartite graph [13]. In their projected networks, the association between two diseases is established if they share a gene implicated in both conditions. The strength of gene-disease associations was quantified by (linear) Pearson correlation, and the strength of disease-disease associations (i.e., their comorbidity) was defined as the number of genes that were significantly associated with both diseases. Although an important innovation, this methodology presents several limitations that might lead to biases in estimating disease associations. Firstly, the method used only a single-omics layer (i.e., genes) to connect diseases, whereas in reality there are multiple omics layers (e.g., genes, proteins, lipids, etc.) connecting diseases. Secondly, simply counting the number of shared intermediate biomarkers is a rather coarse measure that weighs each biomarker equally, irrespective of the the strength of their (Pearson) correlations. Thirdly, the use of Pearson correlation overlooks nonlinear correlations [14]. Lastly, the proposed method does not allow incorporation of higher-order correlations into the projection.
@@
Several studies have sought to address these biases by expanding on the original methodology. For instance, Park et al. implemented cellular networks, including genes and proteins, into their analyses to provide a multi-omics view [15]. To tackle the second limitation, Zhou et al. applied the cosine similarity of disease symptom vectors to quantify disease associations, and defined interaction orders based on the path length between proteins [16]. Taking a different approach, Grosdidier et al. proposed a concept of molecular comorbidity index (MCI) based on proteins common between two diseases to estimate the strength of the association [17]. While these advancements extended the scope to multiple omics or refined the measurement of disease associations, significant methodological gaps remain. Specifically, no existing method provides comprehensive integrative solution that incorporates multiple omics datasets into a single network analysis, employs a more refined quantification of mediating biomarkers, and leverages non-linear correlations for richer expressiveness.
@@
To innovate on existing methodologies and resolve the above gaps, we introduced a multipartite projection method that advances the study of disease networks by: (i) incorporating any number of omics datasets and various diseases; (ii) utilizing the non-parametric Shannon mutual information (MI) correlation; and (iii) employing a weighted approach to count shared intermediate biomarkers and risk factors. MI is a non-parametric metric that captures both linear and nonlinear mutual interdependencies between variables [18], offering a more comprehensive measure of association than conventional methods. By counting biomarkers in a weighted manner, our method additionally provides a finer-grained analyses to quantify the links between diseases. We proposed three increasingly complex and refined definitions for quantifying projected disease links. These definitions not only consider the number of shared biomarkers, but also account for the strength of their nonlinear correlations, and the proximity of their neighboring biomarkers. This approach allows for a more comprehensive understanding of potential shared etiologies compared to direct correlation methods.
@@
We applied the proposed projection method to a dataset from the Young Finns Study (YFS) and selected a projection definition that demonstrated the most effective performance. The dataset encompasses three primary sets of variables: (1) CVD and depression phenotypes (see Table 1 and  2), (2) related risk factors (Table 4), (3) intermediate omics biomarkers (metabolites and lipids). Our method and its resulting weighted multilayer networks allow us to: (i) identify key biomarkers across multiple biological layers that contribute to CVD-depression comorbidity; (ii) assess the relative importance of risk factors by integrating them into the projected network; (iii) identify key biomarkers related to a particular projected link between disease phenotypes or between a risk factor and a disease phenotype. The projection method identified sex and body mass index (BMI) as two important risk factors related to both CVD and depression, and uncovered significant mediating biomarkers underlying their comorbidity. These biomarkers included metabolites such as creatinine, valine, phospholipids in very large high-density lipoproteins (HDL), triglycerides in small low-density lipoproteins (LDL), and apolipoprotein B, as well as lipids such as specific sphingomyelins (SM), phosphatidylcholines (PC), triacylglycerols (TAG) and diacylglycerols (DAG). By accommodating multiple omics datasets and considering non-linear correlations, our method provides a comprehensive, system-level view on the biological pathways and mechanisms contributing to comorbidities. This approach offers a valuable tool for future research, facilitating a deeper understanding of complex disease interactions.
@@ Methods
Our proposed method was applied to the 2007 wave of the Young Finns Study (YFS) [Raitakari2008]. This cross-sectional follow-up study provides a multidimensional dataset that encompasses: (1) CVD-related phenotypes, depressive symptoms, (2) related risk factors (i.e., covariates), (3) intermediate omics biomarkers (metabolites, lipids). In CVD-related variables, the ideal cardiovascular health (CVH) score is a composite measure that typically includes factors such as blood pressure, cholesterol levels, diet, physical activity, and smoking status. Higher scores reflect better health outcomes across these metrics, suggesting a lower risk of cardiovascular disease. The Beck Depression Inventory (BDI) score is a summary variable derived from individual depressive symptoms. A higher BDI score indicates worse mental health, as it reflects a higher level of depressive symptoms. This multidimensional dataset enables the identification of significant biomarkers, and the inclusion of two omics datasets facilitates the use of our multipartite projection approach.
For each participant, the phenotype variables consist of 17 CVD-related indicators and six risk factors (see Table  1 and  4). Depressive variables include the BDI score and 21 individual depressive symptoms contributing to the calculation of the BDI score (see Table 2). The two omics datasets comprise 228 metabolomic variables and 437 lipidomic variables, which were stored in two separate tables. Detailed information about these variables and the methods used for their measurement can be found in Section S1.1 of the Supplementary Information.
@@
We used random sample imputation to replace each missing value of a variable with a random sample drawn from existing values of that variable. This conservative approach minimizes the risk of introducing spurious associations, with the impact of imputation diminishing as the proportion of missing values increases. After imputation, we retained the data points in which all omics data were available, resulting in a final sample of 1686 participants (983 women, 703 men). Continuous variables were discretized by using a quantile-based method, dividing each continuous variable into probabilistically equal-sized buckets. Discretization was needed because part of the data was discrete and information-theoretic metrics require comparable data types. The optimal number of quantiles was determined by Sturges’ rule. Categorical and discrete variables were included without modification.
@@
Before constructing the significant MI network, we first applied a filtering process to minimize double counting and reduce redundant information among variables. The step was necessary because our proposed method incorporates the strength of MI correlations into the projection, and ensuring that each correlation is uniquely represented without redundancy is crucial for the method’s effectiveness. Consequently, some repeated variables and one variable from each highly correlated pair based on MI correlation were excluded from the analysis (see Section S1.1 of the Supplementary Information). These variables included the BDI score, CVD-related variables (such as "fmd40", "fmd40pr", "fmd60", "fmd60pr", "fmd80", "fmd80pr", "maxfmd", "imtmax", "bbmax" and "volscore"), and another 115 variables in metabolites and lipids. This process resulted in a final dataset of 584 variables for further analysis.
@@
The MI can quantify correlations directly between variables of phenotypes, symptoms and risk factors. However, these direct correlations would not be able to suggest any plausible mechanisms (pathways) between them, let alone identify key biomarkers. Our multipartite projection method reconstructs these disease links indirectly as projected scores, based on their MI correlations with intermediate biomarkers.
We provided three definitions for the projected score. In the main analysis, we selected the most effective definition based on a comparative analysis (see Section S2.1 of the Supplementary Information). This definition calculates the projected score as the sum of the average MI correlations between each pair of variables and their shared neighboring nodes (intermediate biomarkers). It is formulated as follows:
@@
This projection definition considers intermediate biomarkers in a weighted manner. When projected to different types of biomarkers, this leads to a weighted, multilayer disease network of CVD phenotypes, depressive symptoms and related risk factors, in which each layer of the network corresponds to the type of biomarker involved in the projection. A stylized representation of this projection definition using average MI correlation is depicted in Figure 1.
@@
The projection method enables us to calculate the contribution of each biomarker (in our case, metabolites and lipids) to the projected score of each link between diseases. Considering the projection definition used in the main analysis and our particular focus on CVD-depression comorbidity, we defined the total contribution score of a biomarker to all possible projected links between CVD phenotypes and depressive symptoms as follows:
@@
where $Y_{k}$ is a biomarker, $X_{i}$ and $X_{j}$ represent pairs of phenotypes and symptoms. Both $f(X_{i};Y_{k})$ and $f(X_{j};Y_{k})$ are non-zero. Such a contribution score enables the identification of significant biomarkers that contribute the most to the overall comorbidity between CVD and depression. For a given pair of $(X_{i},X_{j})$, the method identifies the significant biomarkers that contribute the most to the projected score of the specific link between these diseases.
@@
It is widely recognized that CVD and depression are often linked to behavioral and demographic risk factors [19, 12]. Therefore, risk factors were included in the projection. This approach allows for the assessment of their relative importance concerning specific phenotypes. The relative importance of a risk factor with respect to a particular phenotype is defined as the ratio of the projected score between the phenotype and the risk factor to the total projected score between the phenotype and all risk factors. This can be formulated as:
@@
where $X_{j}$ is a phenotype and $Z_{i}$ is one of the risk factors that has a link to the phenotype. $w(Z_{i};X_{j})$ is the projected score between the phenotype and the risk factor. The relative importance of risk factors considering CVD and depression is computed as the average of $r(Z_{i},X_{j})$ over all CVD-related phenotype and depressive symptoms, respectively. For a particular strong projected link between a risk factor and a disease, we can also identify the significant biomarkers contributing to the projected score using Equation  3.
@@ Results
Significant tripartite MI network and projected multilayer disease networks
@@
The projection method builds on the construction of a significant multipartite MI correlation network. As shown in Figure 2A, this network is tripartite, in which variables are partitioned into three groups: metabolome (blue), lipidome (green), and the combined group of phenotypes/symptoms of CVD (yellow) and depression (red) along with risk factors (purple). It shows only the links between metabolites/lipids and variables in the combined group. The edge width is determined by the mean MI value, which was calculated over 20 iterations of random sample imputation. Node size is proportional to the node’s weighted degree, reflecting the strengths of its connections. To enhance readability and visualization, nodes in the tripartite network were filtered based on their weighted degree, retaining 105 nodes and 669 edges. Consequently, no depressive symptom variables are visible in this network due to their weak correlations with intermediate biomarkers. However, it is important to note that this filtering process was applied only for visualization purposes; the projection analysis itself was performed on the unfiltered network, and depressive symptoms were thus included in the analysis.
@@
Within the tripartite network, certain risk factors such as sex and BMI, CVD-related phenotypes like the ideal CVH score, average diastolic, and systolic blood pressure, exhibit a high weighted degree, indicating extensive correlations with intermediate biomarkers. However, interpreting these correlations based solely on this network poses challenges, as it remains unclear which specific biomarkers are driving these strong associations. This limitation underscores the difficulty in elucidating the underlying causative mechanisms and biological pathways from the correlation network alone.
@@
Figure 2B and C illustrate the metabolomic layer and lipidomic layer of the projected network of CVD-related phenotypes and depressive symptoms, respectively. In these projected networks, the link (edge) width is quantified by the mean projected score over 20 runs of random sample imputation. The node size is proportional to its weighted degree. We only include the projected links between CVD and depression variables due to our focus on the comorbidity between these two conditions. The projected networks are thus displayed in a bipartite layout. In the metabolomic layer of the projected network, we find strong projected links between five individual depressive symptoms (pessimism (’b2’), guilty feelings (’b5’), crying (’b10’), changes in appetite (’b18’), loss of libido (’b21’)) and three CVD-related variables (average diastolic blood pressure (’dkv’), average systolic blood pressure (’syst’), and ideal CVH score (’idealCVH’)). In the lipidomic layer, strong projected links are observed between three depressive symptoms (’b10’, ’b18’ and ’b21’) and the same three CVD-related phenotypes. These strong projected links suggest a high risk of co-occurrence between these phenotypes and symptoms, predicting multimorbid phenotypes of CVD and depression.
@@
To investigate the biological pathways underlying CVD-depression comorbidities, we calculated the mean total contribution score for each metabolite and lipid across all projected links between CVD-related phenotypes and depressive symptoms over 20 runs of random imputation. Figure 4A and B display the top ten mediating metabolites and lipids that contribute the most to these projected CVD-depression links. The top mediating metabolites include creatinine, valine, leucine, phospholipids in very large HDL and small VLDL, triglycerides in small LDL and very small VLDL, and apolipoprotein B. The top mediating lipids consist of several specific sphingomyelins (SM), phosphatidylcholines (PC), triacylglycerols (TAG), diacylglycerols (DAG), phosphatidylglycerols (PG), and phosphatidylethanolamines (PE). These metabolites and lipids were identified as significant biomarkers that link CVD and depression through biological pathways such as inflammation, cholesterol regulation, and atherosclerosis. Most of these metabolites and some lipids are already known, while others, such as SM and PG, are potentially novel biomarkers. Detailed rankings and further information about these biomarkers can be found in Table 3.
@@
Including risk factors in the projection results in a projected multilayer network, in which nodes are partitioned into three groups: risk factors, CVD-related phenotypes, and depressive symptoms (see Figure 5A and B). Observing both metabolomic and lipidomic layers of the projected network, the high weighted degree of sex and BMI indicates their extensive associations with both CVD and depression compared to other risk factors. The relative importance of all risk factors is shown in Table 4. The results suggest that sex and BMI are two important risk factors related to both CVD and depression, which corroborates previous studies [20, 21, 22, 23]. It is important to note that sex can be a covariate or confounder. As a result, we conducted gender-specific analyses (see Section S2.3 of the Supplementary Information). Additionally, age–another well-known risk factor–does not show high importance in the YFS dataset, likely because the YFS participants were all young adults.
We selected two demonstrated high-score projected links ("Sex"–"Appetite (’b18’)" and "Sex"–"idealCVH") to further analyze the contributing biomarkers. In this case, sex may act as a confounding factor, influencing both variables. Due to this confounding effect, it is plausible that appetite and the ideal CVH score show a strong projected score in the projected multilayer networks (see Figure 2 B and C). The significant mediating biomarkers based on the mean total contribution score are shown in Figure 5C, D. Details of these biomarkers can be found in Table S6 and S7. For the link between sex and appetite, significant metabolites include creatinine, free cholesterol in large HDL, and phospholipids, cholesterol and triglycerides in very large VLDL (very low-density lipoprotein). Significant lipids are mainly sphingomyelins (SM) 32:2 and lysophosphatidylcholines (LPC). For the link between sex and the ideal CVH score, significant metabolites encompass phospholipids, free and total cholesterol in large HDL and LDL, apolipoprotein B, and Leucine. Significant lipids are primarily triacylglycerols (TAG) and diacylglycerols (DAG). These identified biomarkers shed light on how risk factors influence the development of CVD and depression.